OpenStudy (anonymous):

Which equation shows Euler's formula? V + E = F + 2 V + E = F - 2 V + F = E + 2 V + F = E - 2

OpenStudy (anonymous):

V - E + F = 2 = V + F = E + 2

OpenStudy (anonymous):

thanks